A 450 kg satellite was launched by a popular TV station to orbit Earth.




How far should it be from the Earth’s surface to experience 4.25 x 103 N




of gravitational force?

1
Expert's answer
2022-03-02T10:14:07-0500
"F=G\frac{Mm}{(R+h)^2}"

"h=\sqrt{\frac{GMm}{F}}-R\\
=\sqrt{\frac{6.67*10^{-11}*5.97*10^{24}*450}{4.25*10^3}}-6378*10^3\\
=1.15*10^5\:\rm m"